# Deep Geometric Moments Promote Shape Consistency in Text-to-3D Generation

### v1

@@ Abstract
To address the data scarcity associated with 3D assets, 2D-lifting techniques such as Score Distillation Sampling (SDS) have become a widely adopted practice in text-to-3D generation pipelines. However, the diffusion models used in these techniques are prone to viewpoint bias and thus lead to geometric inconsistencies such as the Janus problem. To counter this, we introduce MT3D, a text-to-3D generative model that leverages a high-fidelity 3D object to overcome viewpoint bias and explicitly infuse geometric understanding into the generation pipeline. Firstly, we employ depth maps derived from a high-quality 3D model as control signals to guarantee that the generated 2D images preserve the fundamental shape and structure, thereby reducing the inherent viewpoint bias. Next, we utilize deep geometric moments to ensure geometric consistency in the 3D representation explicitly. By incorporating geometric details from a 3D asset, MT3D enables the creation of diverse and geometrically consistent objects, thereby improving the quality and usability of our 3D representations.
@@ 1 Introduction
ControlNet mitigates viewpoint bias to a certain extent, but it performs poorly on views that are significantly scarce in the training dataset. For example, as shown in Figure 4, ControlNet generates a poor-quality image when conditioned on a depth map of a more challenging bottom view. In this paper, we employ deep geometric moments (DGM) [44, 45] to explicitly infuse the geometric properties of a high-fidelity 3D asset into the generated 3D representation, ensuring that even underrepresented views exhibit coherent geometry. Geometric moments define various shape characteristics and can be visualized as projections of the image onto chosen basis functions. Historically, the theoretical development of moments has often relied on specific selections of these bases [47, 17]. Moments have been extensively studied in the vision community for a wide range of applications such as invariant pattern recognition [17, 21, 1, 14], segmentation [35, 15], and 3D shape recognition [48, 39, 12].
@@ 3 Preliminaries
Variational Score Distillation (VSD) [50] treats the 3D representation as a distribution range instead of a single point, as seen in Score Distillation Sampling (SDS). This approach enhances both the quality and diversity of the outputs. VSD utilizes a particle-based update method through the Wasserstein gradient flow. The VSD gradient is formulated as,
@@ 4 Methods
Our objective is to develop high-fidelity, diverse, and geometrically consistent 3D objects from textual descriptions. While existing 2D-to-3D lifting methods can create novel and diverse scenarios based on text inputs, they often face issues with geometric inconsistencies, such as the multi-faced Janus problem, due to a lack of explicit 3D geometric knowledge. In this paper, we aim to close this gap by embedding explicit geometric information into our model using a high-quality 3D image, thus improving the realism of our generated 3D objects. Specifically, upon receiving a text prompt, we identify a suitable high-fidelity 3D image (Section 4.1), and utilize this image to enhance our model’s geometric understanding.
In this paper, we adopt 3D Gaussians for our 3D representations, valuing their flexibility in incorporating geometric priors and their efficiency in rendering. Following [5, 46, 18, 8], our methodology also includes a dual-phase training strategy. The first phase focuses on refining the geometric aspects of the 3D representation (Section 4.2), and the subsequent phase concentrates on enhancing the textural details (Section 4.3).
@@ 4.1 Retrieving high-fidelity 3D object
Upon receiving a text prompt, we select a high-quality 3D object from Objaverse [11], a moderately-sized dataset of diverse 3D assets, to steer the generation process. For every object within Objaverse, we create a corresponding caption using Cap3D [28]. These captions, along with the input text prompt, are then processed through SBert [36] to calculate cosine similarities. The object that shows the highest similarity to the text prompt is chosen for the next steps in the generation sequence.
@@ 4.2 Geometric Refinement
In SDS optimization, we aim to reduce the distribution discrepancy between images rendered from the 3D model and those generated by a pre-trained 2D diffusion model. However, a significant challenge arises from the viewpoint bias inherent in these diffusion models [43, 23, 25], which may hinder the model’s ability to generate diverse views corresponding to different camera settings. This often leads to uniform appearances across various camera angles, contributing to the Janus problem. GSGEN [8] addressed these geometric inconsistencies by incorporating a 3D SDS loss alongside a pre-trained point cloud diffusion model. They found that 3D Gaussians could be effectively adjusted using point cloud priors, a strategy not viable with NeRFs due to their implicit nature. Therefore, to enhance the 3D generation process, in addition to the standard SDS, they also employ Point-E [32], a text-to-point cloud diffusion model, guided by the following equation:
@@
where, $p_{t}$ is noisy Gaussian position, and $\omega_{*}$ and $\epsilon_{*}$ refer to the corresponding weighting function and Gaussian noise, respectively. Applying SDS with Point-E generally results in better geometry compared to standard SDS. However, they still suffer from geometric inconsistencies due to two primary factors. Firstly, the 2D diffusion model suffers from viewpoint bias. Secondly, there is no explicit optimization focused on the overall shape or structure of the 3D representation. We address these challenges in Section 4.2.1 and 4.2.2.
@@ 4.2.1 Controlling 3D Generation
In SDS, the goal is to minimize the distribution gap between images rendered from diffusion models and those from the 3D representation. However, these models typically exhibit a bias towards certain viewpoints, as shown in Figure 2. To address this, we employ ControlNet [56] to counteract this bias. Figure 4 demonstrates that ControlNet, when provided with a conditional image, can generate images from a variety of viewpoints, effectively reducing the viewpoint bias associated with standard diffusion models. In our approach, we utilize a high-fidelity 3D object ($\theta_{h}$) as a control signal for ControlNet. During SDS, for each view rendered from the 3D representation $\theta$, we also project a corresponding view from $\theta_{h}$ using identical camera parameters. We use the depth map of the rendered view as a guiding signal for ControlNet, ensuring that the generated 2D image preserves the essential shape and structure. Although substituting the diffusion model within SDS with ControlNet improves geometric fidelity, SDS continues to demonstrate limited diversity. To address this, we utilize Variational Score Distillation (VSD), which broadens the range of generation for a given prompt. In VSD, a LoRA model conditioned on the camera parameter $c$ is utilized to estimate the noise in the current Gaussian scene, denoted as $\hat{\epsilon}_{\theta}(r_{t},t,c,y)$. Optimizing LoRA with explicit conditioning on camera parameters proves challenging. Therefore, we refine the model to implicitly incorporate camera parameters by substituting $c$ with depth maps ($\theta_{hc}^{d}$). Consequently, LoRA is conditioned on depth maps derived from the high-fidelity images ($\theta_{h}$) corresponding to the camera parameters $c$. The training objective for LoRA is the same as Eq. (3). However, initially, LoRA lacks practical significance and tends to predict random noise. To address this, we initially set the weights of the LoRA term to zero, gradually increasing them over iterations using a cosine schedule. As a result, $\hat{\epsilon}_{\phi}(r_{t},t,y)-\hat{\epsilon}_{\theta}(r_{t},t,c,y)$ term in Eq. (2) is replaced by
@@
where $\hat{\epsilon}_{\phi}(r_{t},t,\theta_{hc}^{d},y)$ is the noise predicted by depth-conditioned ControlNet and $\lambda_{lora}$ is a weighting term. In summary, at each optimization step, we render an RGB image $r$ from $\theta$ and a depth map $\theta_{hc}^{d}$ from $\theta_{h}$ using the same camera parameters $c$. The depth map serves as a conditional image for ControlNet and LoRA. Thus, the optimization step in Eq. (5) is revised as follows:
@@ 4.2.2 Deep Geometric moment
In the section above, we address the viewpoint bias prevalent in diffusion models. ControlNet effectively mitigates this bias to a degree, but it performs poorly on views that are significantly underrepresented in the training data. For instance, as illustrated in Figure 4, ControlNet successfully generates left, right, and front views. However, it fails to produce the more challenging bottom view. In this section, we utilize geometric moments to explicitly induce the geometric properties of a high-fidelity 3D image into the generated 3D representation, ensuring that even underrepresented views maintain coherent geometry.
Classical geometric moments are highly effective at capturing shape information and providing discriminative cues. However, their discriminative power is relatively limited and typically requires the presence of a salient object against a homogeneous background. In response to this limitation, recent work [44, 45] has introduced deep geometric moments (DGM). This approach trains a neural network to learn higher-order geometric moments in an end-to-end differentiable manner. DGM excels at capturing the geometric shapes and structures within 2D images. Figure 4 displays features extracted using the ImageNet pre-trained DGM model.
@@ 4.3 Texture Refinement
The texture refinement stage builds upon geometric refinement, where we progressively enhance and increase the density of the 3D Gaussians. Following the approach detailed in [20], our strategy targets Gaussians for splitting based on their significant view-space spatial gradients. We also incorporate a strategy for densification based on compactness, as suggested by [8], by inserting new Gaussians where the spatial gap between existing ones surpasses their combined radii; the radius of the new Gaussian is adjusted to bridge the remaining gap. Moreover, we periodically remove Gaussians, considering their opacity and size. The loss function applied in this process is identical to geometric refinement (Eq. (8)).
@@ 5.1 Text-to-3D Generation
We assess MT3D’s performance in the text-to-3D generation domain through both qualitative and quantitative comparisons against various state-of-the-art methods [24, 5, 58, 50, 8]. Another notable work, Dreamcontrol [18], also achieves state-of-the-art results; however, we are unable to compare with them as their code is not available at the time of writing this paper. As depicted in Figures 5 and 1, MT3D consistently produces high-quality, well-structured 3D assets with consistent shape and structure. In contrast, alternative approaches often demonstrate significant geometric inconsistencies when responding to the same text prompts. For instance, with the ceramic lion prompt, both Magic3D and Fantasia3D create assets with multiple faces. When prompted with the dog, HiFA generates images with extra eyes near its ears, and Fantasia3D’s model produces a disproportionately large head compared to the body when viewed from the front. Conversely, MT3D consistently delivers accurate and precise geometric renderings of both lions and dogs. Similarly, for challenging prompts like ‘train’ and ‘ostrich,’ where methods such as ProlificDreamer, GSGen, and HiFA fall short in producing photorealistic images and often struggle with multiple facial features, MT3D excels by generating assets that are not only photorealistic but also accurately depict shape and structure.
@@
Figure 6: The results of an ablation study of MT3D.
@@ 5.2 Ablation Study
In this section, we present an ablation study to evaluate the impact of various components in our 3D generation pipeline, specifically focusing on controlling 3D generation, deep geometric moment and the use of a high-fidelity 3D object. Figure 6 showcases the comparison among five generative pipelines. The ‘SDS’ configuration refers to models optimized using Eq. 5 during Stage 1. ‘VSD’ represents model optimized by using VSD (2) instead of SDS (1) in Eq. 5. ‘VSD + $\mathcal{L}_{\text{control}}$’ indicates models trained with the additional Eq. 7. ‘VSD + $\mathcal{L}_{\text{moment}}$’ signifies models trained with both Eq. 5 and the DGM loss ($\mathcal{L}_{\text{dgm}}$). Lastly, ‘VSD + $\mathcal{L}_{\text{control}}$ + $\mathcal{L}_{\text{moment}}$’ describes our proposed pipeline, as detailed in Eq. 8.
@@
Models optimized with the SDS and VSD configuration exhibit various geometric inconsistencies, including multiple faces, indistinct legs, and random noise at the bottom. However, integrating each component of MT3D leads to significant enhancements in overall shape and structure. Although $\mathcal{L}_{\text{dgm}}$ and $\mathcal{L}_{\text{control}}$ appear similar at first glance, notable differences emerge upon closer inspection. For example, the side view of a dog’s leg is better shaped under $\mathcal{L}_{\text{control}}$ than under $\mathcal{L}_{\text{dgm}}$. Conversely, the bottom view is more accurately rendered with $\mathcal{L}_{\text{dgm}}$ than with $\mathcal{L}_{\text{control}}$. This can be attributed to the likelihood that ControlNet’s training dataset predominantly comprises side views of dogs, rather than bottom views, resulting in a better generation of side views compared to bottom views. MT3D leverages both $\mathcal{L}_{\text{dgm}}$ and $\mathcal{L}_{\text{control}}$, resulting in superior representations of the feet, tail, and bottom view, thereby enhancing the overall shape and structure.
Analogous geometric guidance. Instead of using a high-fidelity 3D object that directly matches the text prompt, we explore using 3D objects from a similar category for guidance. For instance, in response to a ‘furry dog’ prompt, we provide guidance using 3D models of a cat and a lion, rather than a dog. As shown in Figure 7, despite leveraging comparable objects from analogous classes, MT3D successfully produces well-structured and high-fidelity 3D representations. This underscores the model’s capability to generalize and effectively utilize geometric cues from similar objects, even in the absence of a precise match.
Figure 7: Illustration of 3D objects generated by our model using various high-fidelity objects corresponding to input text prompt ‘A high quality photo of a furry dog’. Top row represents input high-fidelity object and bottom row represents generated 3D asset.
Low-fidelity 3D object. To assess our model’s generative capabilities, we evaluate its performance using a low-fidelity 3D object. Rather than employing a high-fidelity object, we create a 3D model using the Point-E [32] and then covert it into a mesh object. As depicted in Figure 8, despite limited geometric guidance, our model effectively generates a decent 3D structure with minor geometric inconsistencies. For example, the corgi generated exhibits multiple ears, yet it represents a significant improvement over the representations generated without any guidance, as shown in Figure 6.
Figure 8: Illustration of 3D objects generated by our model using low-fidelity objects corresponding to input text prompt (a) ‘A car made up of cheese’ and (b) ‘A high quality photo of a furry corgi’.

### v2

@@ Abstract
To address the data scarcity associated with 3D assets, 2D-lifting techniques such as Score Distillation Sampling (SDS) have become a widely adopted practice in text-to-3D generation pipelines. However, the diffusion models used in these techniques are prone to viewpoint bias and thus lead to geometric inconsistencies such as the Janus problem. To counter this, we introduce MT3D, a text-to-3D generative model that leverages a high-fidelity 3D object to overcome viewpoint bias and explicitly infuse geometric understanding into the generation pipeline. Firstly, we employ depth maps derived from a high-quality 3D model as control signals to guarantee that the generated 2D images preserve the fundamental shape and structure, thereby reducing the inherent viewpoint bias. Next, we utilize deep geometric moments to ensure geometric consistency in the 3D representation explicitly. By incorporating geometric details from a 3D asset, MT3D enables the creation of diverse and geometrically consistent objects, thereby improving the quality and usability of our 3D representations. Project page and code: https://moment-3d.github.io/
@@ 1 Introduction
ControlNet mitigates viewpoint bias to a certain extent, but it performs poorly on views that are significantly scarce in the training dataset. For example, as shown in Figure 4, ControlNet generates a poor-quality image when conditioned on a depth map of a more challenging bottom view. In this paper, we employ deep geometric moments (DGM)  [33, 46, 47] to explicitly infuse the geometric properties of a high-fidelity 3D asset into the generated 3D representation, ensuring that even underrepresented views exhibit coherent geometry. Geometric moments define various shape characteristics and can be visualized as projections of the image onto chosen basis functions. Historically, the theoretical development of moments has often relied on specific selections of these bases  [49, 18]. Moments have been extensively studied in the vision community for a wide range of applications such as invariant pattern recognition  [18, 22, 1, 14], segmentation  [37, 15], and 3D shape recognition  [50, 41, 12].
@@ 4 Methods
Our objective is to develop high-fidelity, diverse, and geometrically consistent 3D objects from textual descriptions. While existing 2D-to-3D lifting methods can create novel and diverse scenarios based on text inputs, they often face issues with geometric inconsistencies, such as the multi-faced Janus problem, due to a lack of explicit 3D geometric knowledge. In this paper, we aim to close this gap by embedding explicit geometric information into our model using a high-quality 3D object, thus improving the realism of our generated 3D objects. Specifically, upon receiving a text prompt, we identify a suitable high-fidelity 3D object (Section 4.1), and utilize this representation to enhance our model’s geometric understanding.
We adopt 3D Gaussians for our 3D representations, valuing their flexibility in incorporating geometric priors and their efficiency in rendering. Following  [5, 48, 19, 8], our methodology also includes a dual-phase training strategy. The first phase focuses on refining the geometric aspects of the 3D representation (Section 4.2), and the subsequent phase concentrates on enhancing the textural details (Section 4.3).
@@ 4.1 Retrieving high-fidelity 3D object
Upon receiving a text prompt, we select a high-quality 3D object from Objaverse  [11], a moderately-sized dataset of diverse 3D assets, to steer the generation process. For every object within Objaverse, we create a corresponding caption using Cap3D  [29]. These captions, along with the input text prompt, are then processed through SBert  [38] to calculate cosine similarities. The object that shows the highest similarity to the text prompt is chosen for the next steps in the generation sequence. It is important to note that our generation pipeline is compatible with any 3D representation. In this section, we have automated the selection of high-fidelity 3D objects for ease of implementation.
@@ 4.2 Geometric Refinement
In SDS optimization, we aim to reduce the distribution discrepancy between images rendered from the 3D model and those generated by a pre-trained 2D diffusion model. However, a significant challenge arises from the viewpoint bias inherent in these diffusion models  [45, 24, 26], which may hinder the model’s ability to generate diverse views corresponding to different camera settings. This often leads to uniform appearances across various camera angles, contributing to the Janus problem (Figure 2(a)). GSGEN  [8] addressed these geometric inconsistencies by incorporating a 3D SDS loss alongside a pre-trained point cloud diffusion model. They found that 3D Gaussians could be effectively adjusted using point cloud priors, a strategy not viable with NeRFs due to their implicit nature. Therefore, to enhance the 3D generation process, in addition to the standard SDS, they also employ Point-E  [34], a text-to-point cloud diffusion model, guided by the following equation:
@@
where $p_{t}$ is noisy Gaussian position, and $\omega_{*}$ and $\epsilon_{*}$ refer to the corresponding weighting function and Gaussian noise, respectively. Applying SDS with Point-E generally results in better geometry compared to standard SDS. However, they still suffer from geometric inconsistencies due to two primary factors. Firstly, the 2D diffusion model suffers from viewpoint bias. Secondly, there is no explicit optimization focused on the overall shape or structure of the 3D representation. We address these challenges in Section 4.2.1 and 4.2.2.
@@ 4.2.1 Controlling 3D Generation
In this section, we aim to address the Janus problem associated with SDS by mitigating viewpoint bias. Additionally, SDS demonstrates limited diversity as it treats 3D representations as a single point in real space. To overcome these challenges, we incorporate ControlNet  [58] and Variational Score Distillation (VSD)  [52].
Utilizing ControlNet for mitigating viewpoint bias. In our approach, we utilize a high-fidelity 3D object ($\theta_{h}$) as a control signal for ControlNet. Figure 4 demonstrates that ControlNet, when provided with a conditional image, can generate images from a variety of viewpoints, effectively reducing the viewpoint bias associated with standard diffusion models. During SDS, for each view rendered from the 3D representation ($\theta$), we also project a corresponding view from $\theta_{h}$ using identical camera parameters. We use the depth map of the rendered view as a guiding signal for ControlNet, ensuring that the generated 2D image preserves the essential shape and structure.
Improving diversity using VSD. VSD treats the 3D representation as a distribution range instead of a single point, as seen in SDS. This approach enhances both the quality and diversity of the outputs. VSD utilizes a particle-based update method through the Wasserstein gradient flow. The VSD gradient is formulated as,
In our approach, we replace SDS with VSD to broaden the range of generation for a given prompt. However, optimizing LoRA with explicit conditioning on camera parameters proves challenging. Therefore, we refine the model to implicitly incorporate camera parameters by substituting $c$ with depth maps ($\theta_{hc}^{d}$). Consequently, LoRA is conditioned on depth maps derived from the high-fidelity images ($\theta_{h}$) corresponding to the camera parameters $c$. The training objective for LoRA is the same as Eq. (5). However, initially, LoRA lacks practical significance and tends to predict random noise. To address this, we initially set the weights of the LoRA term to zero and gradually increase them over iterations using a cosine schedule. As a result, $\hat{\epsilon}_{\phi}(r_{t},t,y)-\hat{\epsilon}_{\theta}(r_{t},t,c,y)$ term in Eq. (4) is replaced by
@@
where $\hat{\epsilon}_{\phi}(r_{t},t,\theta_{hc}^{d},y)$ is the noise predicted by depth-conditioned ControlNet and $\lambda_{lora}$ is a weighting term.
In summary, as illustrated in Figure 3, at each optimization step, we render an RGB image $r$ from $\theta$ and a depth map ($\theta_{hc}^{d}$) from $\theta_{h}$ using the same camera parameters $c$. The depth map serves as a conditional image for ControlNet and LoRA. Thus, the optimization step in Eq. (3) is revised as follows:
@@ 4.2.2 Deep Geometric Moment
In the previous section, we address the viewpoint bias prevalent in diffusion models. ControlNet effectively mitigates this bias to a degree, but it performs poorly on views that are significantly underrepresented in the training data. For instance, as illustrated in Figure 4, while ControlNet successfully generates left, right, and front views, it fails to produce the more challenging bottom view. In this section, we utilize geometric moments to explicitly induce the geometric properties of a high-fidelity 3D object into the generated 3D representation, ensuring that even underrepresented views maintain coherent geometry.
Promoting shape awareness. Classical geometric moments are highly effective at capturing shape information and providing discriminative cues. However, their discriminative power is relatively limited and typically requires the presence of a salient object against a homogeneous background. In response to this limitation, recent work  [33, 46, 47] has introduced deep geometric moments (DGM). This approach trains a neural network to learn higher-order geometric moments in an end-to-end differentiable manner. DGM excels at capturing the geometric shapes and structures within 2D images. Figure 4 displays features extracted using the ImageNet pre-trained DGM model. In this paper, we minimize the gap between the DGM features obtained from our 3D representations and those of the high-fidelity 3D object to explicitly promote shape awareness.
@@ 4.3 Texture Refinement
The texture refinement stage builds upon geometric refinement, where we progressively refine and increase the density of the 3D Gaussians. Following the approach detailed in  [21], our strategy targets Gaussians for splitting based on their significant view-space spatial gradients. We also incorporate a strategy for densification based on compactness, as suggested by  [8], by inserting new Gaussians where the spatial gap between existing ones surpasses their combined radii; the radius of the new Gaussian is adjusted to bridge the remaining gap. Moreover, we periodically remove Gaussians, considering their opacity and size. The loss function applied in this process is identical to geometric refinement (Eq. (8)).
@@ 5.1 Text-to-3D Generation
We assess MT3D’s performance in the text-to-3D generation domain through both qualitative and quantitative comparisons against various state-of-the-art methods  [25, 5, 60, 52, 8]. Another notable work, Dreamcontrol  [19], also achieves state-of-the-art results; however, we are unable to compare with them as their code is not available at the time of writing this paper. As depicted in Figures 5 and 1, MT3D consistently produces high-quality, well-structured 3D assets with consistent shape and structure. In contrast, alternative approaches often demonstrate significant geometric inconsistencies when responding to the same text prompts. For instance, with the ceramic lion prompt, both Magic3D and Fantasia3D create assets with multiple faces. When prompted with the dog, HiFA generates images with extra legs, and Fantasia3D’s model outputs a dog with disoriented legs and body. Conversely, MT3D consistently delivers accurate and precise geometric renderings of both lions and dogs. Similarly, for challenging prompts like ‘train’ and ‘ostrich,’ where methods such as ProlificDreamer, GSGen, and HiFA fall short in producing photorealistic images and often struggle with multiple facial features, MT3D excels by generating assets that are not only photorealistic but also accurately depict shape and structure.
@@
Figure 6: Results of an ablation study on MT3D with the text prompt: "A high-quality photo of a furry dog." The reference object represents a high-fidelity 3D model used to guide MT3D’s generation process. Additional ablation studies can be found in Figure 10 of the appendix.
@@ 5.2 Ablation Study
In this section, we present an ablation study to evaluate the impact of various components in our 3D generation pipeline, specifically focusing on controlling the 3D generation, deep geometric moment and the use of a high-fidelity 3D object. Figure 6 showcases the comparison among five generative pipelines. The ‘SDS’ configuration refers to models optimized using Eq. 3 during Stage 1. ‘VSD’ represents the model optimized by using VSD (4) instead of SDS (1) in Eq. 3. ‘VSD + $\mathcal{L}_{\text{control}}$’ indicates models trained with the additional Eq. 7. ‘VSD + $\mathcal{L}_{\text{moment}}$’ signifies models trained with both Eq. 3 and the DGM loss ($\mathcal{L}_{\text{dgm}}$). Lastly, ‘VSD + $\mathcal{L}_{\text{control}}$ + $\mathcal{L}_{\text{moment}}$’ describes our proposed pipeline, as detailed in Eq. 8.
@@
Models optimized with the SDS and VSD configuration exhibit various geometric inconsistencies, including multiple faces, indistinct legs, and random noise at the bottom. However, integrating each component of MT3D leads to significant enhancements in overall shape and structure. Although $\mathcal{L}_{\text{dgm}}$ and $\mathcal{L}_{\text{control}}$ appear similar at first glance, notable differences emerge upon closer inspection. For example, in the generated representation of a ‘furry dog’, the side view of the dog’s leg is better shaped under $\mathcal{L}_{\text{control}}$ than under $\mathcal{L}_{\text{dgm}}$. Conversely, the bottom view is more accurately rendered with $\mathcal{L}_{\text{dgm}}$ than with $\mathcal{L}_{\text{control}}$. This can be attributed to the likelihood that ControlNet’s training dataset predominantly comprises side views of dogs, rather than bottom views, resulting in a better generation of side views compared to bottom views. Thus, ControlNet focuses more on learning the general aspects of shape, while DGM primarily learns shape and structure from the reference object. By combining $\mathcal{L}_{\text{dgm}}$ and $\mathcal{L}_{\text{control}}$, MT3D delivers superior fidelity across all perspectives. Additional ablation studies in Section A.2 confirm consistent results across various text prompts.
Figure 7: (left) 3D objects generated by MT3D using low-fidelity reference objects from Point-E. (Right) 3D objects generated by MT3D using analogous reference objects. Please refer to Section A.1 of the appendix for more illustrations.
Absence of high-fidelity reference 3D object: MT3D relies on reference objects for geometrically coherent 3D representations. We evaluate its performance using two alternative guidance types in the absence of high-fidelity references. First, we use low-fidelity or geometrically inconsistent 3D representations generated by other text-to-3D models, such as Point-E  [34]. Second, instead of using a high-fidelity 3D object that directly matches the text prompt, we explore using 3D objects from a similar category for guidance. For example, in response to a ‘ceramic lion’ prompt, we use a 3D model of a horse as guidance. As shown in Figure 7, our model demonstrates the ability to generate well-structured and high-fidelity 3D representations, even when the geometric guidance is limited or geometrically inconsistent. We conduct exhaustive experiments with various types of reference objects in Section A.1 of the appendix.